OpenStudy (vortish):

find the least common multiple x^2-11x,x^2

OpenStudy (anonymous):

i can do this to certain extent take the frist term ------- x^2-11x = x(x-11) take the second term------ x^2 = x(x) now write the factor which is common in both these terms first

OpenStudy (vortish):

x is common in both terms

OpenStudy (anonymous):

yes so if we take it out from both what is left

OpenStudy (vortish):

x-11

OpenStudy (anonymous):

multiply x-11 with the common facor

OpenStudy (vortish):

x^2-11x

OpenStudy (anonymous):

i think this is the LCM

OpenStudy (vortish):

ok

OpenStudy (anonymous):

wait

OpenStudy (vortish):

got it right thanks riz

OpenStudy (anonymous):

its x^2(x-11) ok i was confused sorry